Kate is 12 years old. His age is 4 times the age of Robbie. Find the age of Robbie.
Answer by TutorDelphia(193)

We can reason this problem out. If Kate is 12 and 4 times Robbie's age, Robbie must by 3 because 3*4=12

Doing this using algebra we see that:
k=12
k=4*r
12=4*r
r=3